# Agentic AI Reasoning for Mobile Edge General Intelligence: Fundamentals, Approaches, and Directions

### v1

@@ Abstract
The rapid advancement of large language models (LLMs) has enabled an emergence of agentic artificial intelligence (AI) with powerful reasoning and autonomous decision-making capabilities. This integration with edge computing has led to the development of Mobile Edge General Intelligence (MEGI), which brings real-time, privacy-preserving reasoning to the network edge. However, deploying LLM-based agentic AI reasoning in MEGI environments poses significant challenges due to the high computational demands of reasoning and the limited resources of edge devices. To address these challenges, we propose a joint optimization framework for efficient LLM reasoning deployment in MEGI. First, we review methods that enhance LLM reasoning capabilities, such as Chain-of-Thought (CoT) prompting, Supervised Fine-Tuning (SFT), and Mixture of Experts (MoE). Next, we present a distributed framework that addresses two correlated aspects: reasoning enhancement through adaptive CoT prompting and scalable deployment through distributed MoE architecture. The framework dynamically activates expert networks and adjusts reasoning depth based on task complexity and device capabilities. We further conduct experimental evaluations in mobile edge environments. Experimental results demonstrate the framework’s effectiveness in balancing reasoning quality with resource efficiency, validating the practical viability of deploying sophisticated LLM reasoning capabilities in resource-constrained MEGI environments.
@@ I Introduction
Recent advances in artificial intelligence (AI) demonstrate that reasoning is fundamental to intelligence. The recently proposed DeepSeek-R1 model has demonstrated superior performance in mathematics and coding competitions by employing sophisticated reasoning strategies such as self-verification, reflection, and adaptive decision-making [1]. These results underscore the central role of reasoning in enhancing large language models (LLMs). Strengthened reasoning reduces hallucinations, improves performance on complex analytical tasks, and increases decision reliability through systematic verification and error correction. As a result, LLMs equipped with advanced reasoning can handle cognitive tasks such as mathematical problem solving, logical deduction, and multi-step analytical reasoning [2]. Building on this progress, reasoning-enhanced LLMs have driven the development of agentic AI, a paradigm in which LLMs evolve from passive text generation to autonomous, goal-directed behaviour. Agentic AI systems integrate planning, memory, conditional logic, and tool use, where LLM-based reasoning serves as the core of these capabilities. Unlike traditional systems with fixed input-output mappings, agentic AI dynamically decomposes tasks and formulates multi-step strategies for evolving environments [3]. Concurrently, the convergence of agentic AI with edge computing has led to the development of Mobile Edge General Intelligence (MEGI) [4]. It represents an evolution from traditional task-specific edge intelligence toward versatile systems that handle diverse cognitive tasks.
@@
However, deploying LLM-based agentic AI systems with reasoning capability in MEGI environments introduces significant technical challenges[5]. Edge devices operate under tight resource constraints, including limited computational power, memory capacity, and energy budgets. Furthermore, the sophisticated reasoning mechanisms of agentic AI impose substantial computational demands. The dynamic nature of mobile environments introduces additional complexity through variable network conditions, device mobility, and heterogeneous hardware capabilities. Existing approaches often struggle to jointly optimize for these constraints while preserving the quality and efficiency of LLM reasoning [6, 7, 8, 9]. Recent studies further highlight the severity of this limitation: for instance, when GPU resources are constrained, aggressive quantization is often required for edge deployment. However, 3-bit quantized models can suffer accuracy degradations of up to 40% compared with higher-precision variants [7].
@@
To address these challenges, this paper proposes a joint optimization framework that enables efficient deployment of LLM-based agentic AI systems with reasoning in MEGI environments. Our framework enhances reasoning via adaptive Chain-of-Thought (CoT) prompting. It also ensures scalable deployment through a distributed Mixture of Experts (MoE) architecture. CoT prompting enables step-by-step reasoning processes that can be computationally managed. It dynamically adjusts reasoning depth based on task complexity and resource availability. The distributed MoE architecture complements this by enabling sophisticated reasoning processes to be distributed across edge devices, thereby maintaining reasoning quality while achieving computational scalability. Through comprehensive analysis and experimental validation, this work establishes a foundation for reasoning-enabled AI in wireless communication systems. It provides methodologies for implementing and evaluating reasoning mechanisms in resource-constrained environments, and develops optimization strategies for diverse edge computing scenarios. The contributions of this work can be summarized as follows:
@@
  • We provide an overview of methods for enhancing LLM agentic AI reasoning capabilities across different phases of model development, including pre-training (e.g., model scaling and improved model architecture), fine-tuning (e.g., Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF)), and inference (e.g., CoT prompting and self-consistency). This foundational analysis sets the stage for integrating these advanced reasoning into MEGI.
@@
  • We propose a novel joint optimization framework for LLM agentic AI reasoning in MEGI, featuring a distributed MoE architecture integrated with CoT prompting. This framework systematically minimizes total system energy consumption by managing token assignment, transmission power, and the dynamic CoT reasoning depth. Moreover, it meets critical latency, inference quality, and hardware constraints for edge deployment.
@@
  • We present experimental evaluations, including both local deployment validation and system-level performance assessment. The local deployment validation demonstrates the effectiveness of reasoning enhancement techniques. Furthermore, the system-level evaluation validates the joint optimization framework’s ability to balance reasoning quality with resource efficiency. The experimental results provide evidence for the practical viability of deploying sophisticated LLM reasoning capabilities in MEGI environments.
@@ II-B1 Pre-Training Phase
  • Model Scaling: Model scaling refers to increasing model size by expanding parameters, layers, or architecture dimensions. This approach establishes robust foundational reasoning capabilities, with larger models demonstrating emergent reasoning behaviors, particularly in complex multi-step problems requiring extensive knowledge integration [12]. For instance, the progression from GPT-3 (175 billion parameters) to GPT-4 (estimated 1.76 trillion parameters) demonstrates significant improvements in reasoning performance. Empirical benchmarks, such as the GSM8K mathematical reasoning dataset, show GPT-4 achieving 92% accuracy, a substantial increase from GPT-3.5’s 57% [13].
  • New Model Architectures: Emerging designs such as Mixture of Experts (MoE) and Multi-Agent systems aim to enhance reasoning while maintaining computational efficiency. MoE frameworks utilize sparse activation patterns where only specialized expert networks process each input, while Multi-Agent architectures enable collaborative reasoning through distributed intelligent agents specializing in different problem aspects. These new model architectures are crucial for balancing reasoning capability with computational efficiency, particularly in resource-constrained environments, enabling parallel processing of different reasoning components while maintaining practical deployment feasibility. For example, Yi et al. [9] developed EdgeMoE, a sparse architecture designed for mobile device deployment. By selectively activating relevant expert networks, it achieves enhanced reasoning with 5% to 18% reduced memory usage and 1.2 to 2.7 times faster inference, enabling real-time performance for models exceeding 10 billion parameters.
@@ II-B3 Inference Phase
Overall, these methods demonstrate that reasoning performance can be improved across different stages of LLM development. Pre-training methods expand the model’s intrinsic reasoning capacity, fine-tuning approaches align reasoning with human preferences, and inference-time techniques make reasoning more explicit and controllable. However, these methods are typically studied in isolation and assume sufficient computational resources. When deployed in MEGI environments, their effectiveness may diminish due to energy, latency, and hardware limitations. Therefore, a joint optimization framework is essential to sustain reasoning quality while ensuring practical efficiency in edge deployment.
@@ III-A2 Design Principles
  • On-Demand Expert Activation: A lightweight gating network selectively activates only the most relevant subnetworks for each input token, reducing unnecessary computation and enabling efficient scaling across distributed edge devices. In the system workflow, this is realized through the BS Control Unit’s gating network, which assigns tokens to top-$k$ experts based on relevance scores.
  • Joint Communication Computation Awareness: Task scheduling simultaneously considers device compute capabilities, channel quality, transmission power, and application-specific latency and accuracy requirements, enabling end-to-end optimization. This principle is implemented via an optimization module, which guides CoT step selection, token distribution, and power control.
  • Hierarchical Collaboration: The BS manages global scheduling and result aggregation, while edge devices execute expert networks, establishing a clear separation between control and execution. This collaboration is reflected in the modular organization of the BS Control Unit and distributed edge experts.
@@ III-A3 System Components and Workflow
As illustrated in Fig. 2, our proposed framework enables efficient and scalable LLM inference in MEGI environments. It comprises three components: the BS Control Unit, distributed Edge Devices (Experts), and the integrated CoT Reasoning Module. These components collaboratively execute an optimized, distributed inference workflow tailored for resource-constrained and latency-sensitive settings.
@@
The BS Control Unit serves as the central coordinator of the inference process. It initiates the end-to-end pipeline when a user submits a natural language query. The BS first tokenizes the input and transforms each token into high-dimensional embeddings through the model’s initial layers, including self-attention blocks. The BS then invokes the MoE gating network, which computes soft assignment scores for each token across the pool of available expert networks. Based on these scores, the top-$k$ experts for each token are selected to participate in the inference. The BS subsequently batches tokens according to expert assignments and transmits them to the selected edge devices via downlink channels. Transmission is power-adaptive to ensure both energy efficiency and communication reliability.
The Edge Devices, each equipped with a dedicated expert network, act as distributed computation nodes that enable system scalability within the MEGI infrastructure. Each expert is a specialized feed-forward neural module derived from a sparsely activated MoE layer in the LLM architecture, with capabilities enhanced by prior SFT on domain-specific or task-specific datasets. This fine-tuning refinement enables each expert to achieve higher baseline accuracy and improved computational efficiency for its specialized inference tasks, as demonstrated in Section II-B2. As a result, the framework benefits from SFT-enhanced experts, leading to better overall performance and more robust CoT reasoning on complex queries. These devices execute expert-specific inference tasks assigned by the BS and operate under constraints of device-local memory, computational resources and communication interfaces. Depending on task complexity and available resources, each device can process multiple tokens simultaneously. Upon receiving token batches, edge devices perform integrity checks using cyclic redundancy codes (CRC).
Each expert device integrates a CoT Reasoning Module to enhance reasoning quality on complex tasks. After performing its assigned feed‑forward pass, the expert prompts itself to generate one or more intermediate reasoning steps, essentially articulating its logical progression before providing the final output. This structured reasoning process allows the expert to decompose multifaceted queries into manageable subproblems. However, each CoT step requires an additional forward pass and produces extra token embedding vectors, which increases computational load and output size [11]. To balance reasoning quality with scalability requirements, our implementation includes a reasoning-depth parameter that can be configured. The expert produces one reasoning step at a time and stops once the configured depth or a task-specific quality threshold is reached. The resulting chain of logical outputs is then encoded and sent back with the final token representations.
After local inference, each edge device transmits the intermediate outputs back to the BS over the uplink. The BS aggregates the returned vectors according to their original gating weights to reconstruct the final token-level outputs. Post-processing steps, such as normalization, decoding, and response formatting, are applied before the final result is delivered to the end user. This hybrid architecture and workflow synergistically combine scalability through distributed MoE and reasoning enhancement through adaptive CoT to support efficient, interpretable, and scalable LLM inference across distributed MEGI environments.
@@ IV-A Local Deployment Validation
To establish a foundational understanding of reasoning quality improvements through SFT and CoT prompting, we evaluated locally deployed Gemma-2B [15] models on a mobile edge device. This validation assesses the impact of these methods on reasoning, guiding our distributed system design. The experiment was performed on a Huawei MateBook 14 with 16GB RAM and Intel Core i7-10510U CPU, representing typical mobile edge device capabilities. We compared two model variants: the foundational Gemma-2B-base model and the fine-tuned Gemma-2B-IT model. These lightweight LLMs can be feasibly deployed at the edge, and the setup can be extended to other models. Each model variant was evaluated under two prompting conditions: direct answer generation and structured CoT reasoning.
Figure 3: Summary of output correctness and inference time for Gemma-2B and Gemma-2B-IT with and without CoT prompts on a mobile edge device.
The results in Figure 3 show that the Gemma-2B-IT model provides more accurate answers than the base model with faster response times. The comparison between base and fine-tuned models demonstrates the importance of model selection in resource-constrained environments. While SFT models require initial training overhead, they provide substantial operational efficiency gains that justify the investment for deployment scenarios with sustained usage patterns. CoT prompting improves interpretability but increases latency, and in the base model sometimes introduces additional errors. The Gemma-2B-IT model with direct prompting achieves an optimal balance of accuracy and latency, while CoT prompting provides detailed reasoning at the expense of slightly longer inference times. These findings underscore the trade-offs between model type and prompting strategy, justifying adaptive selection in distributed deployments.

### v2

@@ Abstract
The rapid advancement of large language models (LLMs) has enabled an emergence of agentic artificial intelligence (AI) with powerful reasoning and autonomous decision-making capabilities. This integration with edge computing has led to the development of Mobile Edge General Intelligence (MEGI), which brings real-time, privacy-preserving reasoning to the network edge. However, deploying LLM-based agentic AI reasoning in MEGI environments poses significant challenges due to the high computational demands of reasoning and the limited resources of edge devices. To address these challenges, we propose a joint optimization framework for efficient LLM reasoning deployment in MEGI. First, we systematically review enhancement methods to identify mechanisms suitable for edge adaptation. Subsequently, we present a distributed framework that synergizes reasoning enhancement via adaptive CoT prompting with scalable deployment through a distributed MoE architecture. An important innovation of this approach involves modeling reasoning depth as a dynamic network resource variable, which is optimized jointly with expert activation and transmission power. This mechanism allows the system to dynamically regulate expert networks and reasoning complexity according to task requirements and device capabilities. Experimental evaluations in mobile edge environments demonstrate that the proposed framework effectively balances reasoning quality and resource efficiency. The results show that with less than one second of additional inference time, both accuracy and latency satisfaction rate can reach 90%, validating the practical viability of deploying sophisticated LLM reasoning in resource-constrained MEGI systems.
@@ I Introduction
Recent advances in artificial intelligence (AI) demonstrate that reasoning is fundamental to intelligence. The recently proposed DeepSeek-R1 model has demonstrated superior performance in mathematics and coding competitions by employing sophisticated reasoning strategies such as self-verification, reflection, and adaptive decision-making [4]. This capability emphasizes a paradigm shift toward Agentic AI, where large language models (LLMs) evolve from passive text generation to autonomous, goal-directed behavior. By integrating planning, memory, and tool usage, Agentic AI dynamically decomposes tasks and formulates multi-step strategies to navigate evolving environments [12, 15]. Deploying autonomous agents leads to an introduction of Mobile Edge General Intelligence (MEGI) [14]. While cloud-centric solutions such as ChatGPT [1] offer substantial computational capacity, they often fail to meet the strict requirements of real-time tasks, including automated fault detection and privacy-sensitive interactions. In such scenarios, the propagation latency and data sovereignty risks associated with cloud offloading become prohibitive. MEGI mitigates these limitations by instantiating reasoning capabilities directly at the network edge, thereby ensuring service continuity and privacy preservation.
@@
However, deploying LLM-based agentic AI systems with reasoning capability in MEGI environments introduces significant technical challenges[13]. Edge devices operate under tight resource constraints, including limited computational power, memory capacity, and energy budgets. Furthermore, the sophisticated reasoning mechanisms of agentic AI impose substantial computational demands. The dynamic nature of mobile environments introduces additional complexity through variable network conditions, device mobility, and heterogeneous hardware capabilities. Existing optimization strategies frequently struggle to reconcile these constraints with the demands of high-quality reasoning [2, 6, 8, 11]. Although recent studies have addressed specific aspects of edge deployment, such as service offloading [2], model quantization [6], or sparse activation architectures like EdgeMoE [11], they typically view the inference process as a static workload. This oversight fails to account for the dynamic characteristics of reasoning-enhanced tasks, where inference quality fluctuates with reasoning depth, and communication overhead scales with intermediate token transmission. Consequently, static methods are ill-suited for scenarios requiring flexible computation allocation among distributed experts. For instance, aggressive quantization is often employed to fit models onto constrained GPUs, yet this approach can degrade accuracy by up to 40% compared to higher-precision variants [6].
@@
To address these challenges, this paper proposes a joint optimization framework that enables efficient deployment of LLM-based agentic AI systems with reasoning in MEGI environments. Rather than simply combining existing techniques, our framework establishes a theoretical link between reasoning quality and network resources. The framework enhances reasoning via adaptive Chain-of-Thought (CoT) prompting and ensures scalability through a distributed Mixture of Experts (MoE) architecture. Specifically, CoT prompting decomposes complex problems into manageable steps, allowing the system to dynamically adjust reasoning depth based on task complexity and resource availability. Complementarily, the distributed MoE architecture allocates tasks across edge devices, maintaining high reasoning quality while achieving computational scalability. The contributions of this work can be summarized as follows:
@@
  • We provide an overview of methods for enhancing LLM agentic AI reasoning capabilities across different phases of model development, including pre-training (e.g., model scaling and improved model architecture), fine-tuning (e.g., Supervised Fine-Tuning (SFT), Reinforcement Learning from Human Feedback (RLHF)), and inference (e.g., CoT prompting and self-consistency). This analysis highlights technical prerequisites for adapting reasoning-capable agents to edge environments.
@@
  • We propose a novel joint optimization framework for LLM agentic AI reasoning in MEGI, featuring a distributed MoE architecture integrated with CoT prompting. A key distinction of our framework is the introduction of a reasoning-aware resource management mechanism. We theoretically model the trade-off between reasoning depth and energy consumption, enabling the systematic minimization of total system energy by jointly optimizing token assignment, transmission power, and adaptive reasoning depth.
@@
  • We conduct extensive experimental evaluations, ranging from local deployment validation to system-level performance assessment. The results verify the effectiveness of the reasoning enhancement techniques and demonstrate the framework’s ability to balance reasoning quality with resource efficiency. These findings provide empirical evidence supporting the practical viability of sophisticated LLM reasoning in MEGI environments.
@@ II-B1 Pre-Training Phase
  • Model Scaling: Model scaling refers to increasing model size by expanding parameters, layers, or architecture dimensions. This approach establishes robust foundational reasoning capabilities, with larger models demonstrating emergent reasoning behaviors, particularly in complex multi-step problems requiring extensive knowledge integration. For instance, the progression from GPT-3 (175 billion parameters) to GPT-4 (estimated 1.76 trillion parameters) demonstrates significant improvements in reasoning performance. Empirical benchmarks, such as the GSM8K mathematical reasoning dataset, show GPT-4 achieving 92% accuracy, a substantial increase from GPT-3.5’s 57% [1].
  • New Model Architectures: Emerging designs such as Mixture of Experts (MoE) and Multi-Agent systems aim to enhance reasoning while maintaining computational efficiency. MoE frameworks utilize sparse activation patterns where only specialized expert networks process each input, while Multi-Agent architectures enable collaborative reasoning through distributed intelligent agents specializing in different problem aspects. These new model architectures are crucial for balancing reasoning capability with computational efficiency, particularly in resource-constrained environments. For example, Yi et al. [11] developed EdgeMoE, a sparse architecture designed for mobile device deployment. By selectively activating relevant expert networks, it achieves enhanced reasoning with 5% to 18% reduced memory usage and 1.2 to 2.7 times faster inference, enabling real-time performance for models exceeding 10 billion parameters.
@@ II-B3 Inference Phase
In summary, Fig. 1 highlights that while existing methods improve reasoning, they exhibit distinct trade-offs: pre-training offers intrinsic capacity but high cost; fine-tuning provides specialization but requires data; and inference techniques enhance precision at the expense of latency. By systematically listing these pros and cons, Fig. 1 highlights that existing methods are typically studied in isolation and assume sufficient computational resources. When deployed in MEGI environments, their effectiveness may diminish due to energy, latency, and hardware limitations. Therefore, a joint optimization framework is essential to sustain reasoning quality while ensuring practical efficiency in edge deployment.
@@ III-A2 Design Principles
  • Unlike conventional token-level routing, our framework evaluates the global semantic relevance between the complete input query and expert profiles. By activating the most suitable edge device to handle the entire inference task, we minimize the communication overhead associated with frequent token exchange and ensure that domain-specific queries, such as coding or mathematics, are routed to experts with matching capabilities [11].
  • Joint Communication Computation Awareness: Task scheduling simultaneously considers device compute capabilities, channel quality, transmission power, and application-specific latency and accuracy requirements, enabling end-to-end optimization. This principle is implemented via an optimization module, which guides CoT step selection, routing principle, and power control.
  • Hierarchical Collaboration: The Base Station (BS) manages global scheduling and result aggregation, while edge devices execute expert networks, establishing a clear separation between control and execution. This collaboration is reflected in the modular organization of the BS Control Unit and distributed edge experts.
@@ III-A3 System Components and Workflow
As illustrated in Fig. 2, the proposed framework comprises three primary components: the BS Control Unit, distributed Edge Experts, and the integrated CoT Reasoning Module. These components execute a collaborative inference workflow optimized for resource-constrained settings.
@@
The BS Control Unit serves as the central coordinator of the inference process. It initiates the end-to-end pipeline when a user submits a natural language query. To implement task-level routing, the BS first analyzes the input query to extract its global feature representation. This representation is fed into the global gating network, which computes a relevance score for each edge expert, quantifying the match between the task’s domain and the experts’ specialized capabilities. Based on these scores, the BS assigns the entire inference task to the top-ranked edge expert. Instead of decomposing the task into tokens for disjoint processing, the BS dispatches the complete query to the selected edge device via downlink channels. This avoids the high communication overhead associated with frequent token exchange. Transmission is power-adaptive to ensure both energy efficiency and communication reliability.
The Edge Devices act as distributed computation nodes that enable system scalability within the MEGI infrastructure. In our framework, experts are implemented as standalone small models. Each expert is hosted on an edge device and is specialized through prior SFT on domain-specific datasets. This fine-tuning refinement enables each expert to achieve higher baseline accuracy and improved computational efficiency for its specialized inference tasks, as in Section II-B2. It is worth noting that while this work focuses on architectural sparsity and adaptive reasoning depth, our framework differs from model compression techniques. The deployed expert models can be readily integrated with post-training quantization to further reduce memory footprint and execution latency without altering the proposed optimization logic.
The CoT Reasoning Module is integrated within each expert to govern the inference process. Once assigned a task, the edge device processes the input independently. To handle complex queries, the system explicitly activates CoT prompting, generating a sequence of intermediate reasoning tokens, or "thoughts," prior to the final answer. The length of this sequence, defined as the reasoning depth, is a decision variable. Increasing reasoning depth allows the model to decompose multifaceted problems into manageable sub-steps, thereby enhancing accuracy. However, this process linearly increases computational load and energy consumption [9]. Consequently, our framework treats reasoning depth as a dynamic resource parameter: the system optimizes the trade-off between the performance gains of deeper reasoning and the associated resource costs, instructing the expert to terminate generation once the optimal depth is reached. Specifically, once the optimal reasoning depth is determined by the BS, it is enforced by constructing a prompt that explicitly instructs the expert to solve the problem following a specified number of reasoning steps. Upon completion, the edge device transmits the final response, including both the reasoning steps and the answer, back to the BS for formatting and delivery.
@@ IV-A Local Deployment Validation
To establish a basic understanding of reasoning quality improvements through SFT and CoT prompting, we evaluated locally deployed Qwen3 [10] models on a mobile edge device. To ensure domain relevance, we used the TeleQnA dataset[7], a specialized benchmark comprised of diverse telecommunications standards and signal processing questions. The experiment was performed on a device with an integrated GPU delivering 2 TFLOPS (FP16) and 16 GB of Unified Memory. This configuration represents the performance baseline for edge devices capable of hosting billion-parameter models under constrained power budgets. We compared two model variants: the foundational Qwen3-0.6B-Base model and the fine-tuned Qwen3-0.6B model. These lightweight LLMs can be feasibly deployed at the edge, and the setup can be extended to other models. Both models were tested under direct answer generation and CoT reasoning modes.
Figure 3: Summary of output correctness and inference time for Qwen3-0.6B and Qwen3-0.6B-Base with and without CoT prompts on a mobile edge device.
The results in Fig. 3 show that the Qwen3-0.6B model provides more accurate answers than the base model with faster response times. The comparison between base and fine-tuned models demonstrates the importance of model selection in resource-constrained environments. While SFT models require initial training overhead, they provide substantial operational efficiency gains that justify the investment for deployment scenarios with sustained usage patterns. CoT prompting improves interpretability but increases latency, and in the base model sometimes introduces additional errors. The Qwen3-0.6B-Base model with direct prompting achieves an optimal balance of accuracy and latency, while CoT prompting provides detailed reasoning at the expense of slightly longer inference times. We performed independent inference runs on TeleQnA tasks [7]. The specialized Qwen3-0.6B model achieves a high accuracy of 49.7% with CoT prompting, significantly outperforming the Base model (28.9%). These empirical observations directly inform the design of our distributed MEGI framework. First, the superior efficiency of SFT models motivates our architectural choice of employing specialized, fine-tuned experts within the MoE layer to ensure high baseline competency under edge resource constraints. Second, the explicit trade-off between CoT-induced latency and reasoning accuracy necessitates the dynamic control mechanism. Rather than applying a static CoT setting, we design the system to treat reasoning depth as an adaptive decision variable—increasing depth to boost accuracy for complex queries when channel conditions permit, while reducing it to meet strict latency deadlines during congestion.